OpenStudy (anonymous):

Does 5a³+15a²-6a-18 factor to (5a²-6)(a+3)?

OpenStudy (anonymous):

A good way to check is to foil out and compare.

OpenStudy (anonymous):

I dont think i foil right

Directrix (directrix):

5a³+15a²-6a-18 = 5a² ( a + 3) - 6 ( a + 3) = (a + 3) ( 5a² - 6). Check the accuracy of this factoring by grouping.

Directrix (directrix):

FOIL -> First, Outer, Inner, Last (5a²-6)(a+3)

Directrix (directrix):

First terms are those that appear first in each set of parentheses. 5a² and a. So what is 5a² times a?

Directrix (directrix):

5a² times a ? @JaneDoe100

OpenStudy (anonymous):

sorry i was making a factoring darwing

Directrix (directrix):

We are learning to FOIL now.

Directrix (directrix):

No.

OpenStudy (mertsj):

@Directrix http://openstudy.com/study#/updates/513e836ee4b029b0182c23ed

OpenStudy (anonymous):

5a³

Directrix (directrix):

5 a^2 means 5 times a times a. Then, multiply that by a. So, what is 5 times a times a times a?

OpenStudy (anonymous):

not 5a³?

Directrix (directrix):

5 times a times a times a = 5a³. (F) O is for the outer terms. The outer terms are on the outside. What does 5A^2 times 3 =

OpenStudy (anonymous):

15a²

Directrix (directrix):

15a² (O) I is for inner. The inner terms are (-6) and (a). What is the product of (-6) and (a) ?

OpenStudy (anonymous):

-6a?

Directrix (directrix):

-6a (I) L is for last. What is the product of (-6) and (3) ?

Directrix (directrix):

@JaneDoe100

OpenStudy (anonymous):

-12

OpenStudy (anonymous):

i think i get what i was doing wrong now thanks!!!

Directrix (directrix):

-18 for L (-6)*(3) = -18

Directrix (directrix):

F) 5a^3 O) 15a^2 I) -6a L) -18

OpenStudy (anonymous):

so its (5a²-6)(a+3) because that factored out to the right problem

Directrix (directrix):

(5a²-6)(a+3) = 5a^3 + 15a^2 -6a -18 = 5a^3 + 15a^2 - 6a - 18.

Directrix (directrix):

Yes.

OpenStudy (anonymous):

thanks